Circular No. 7429

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V1333 AQUILAE R. Jain, C. Bailyn, and P. Coppi, Yale University; M. Garcia, Harvard-

Smithsonian Center for Astrophysics; A. Levine, Massachusetts Institute of Technology; J. Orosz, Utrecht University; P. Lu, Western Connecticut State University; and J. Espinoza and D. Gonzalez, Cerro Tololo Interamerican Observatory (CTIO), report YALO consortium observations: “The optical outburst of V1333 Aql = Aquila X-1, which began on May 11.23 UT (IAUC 7423), did not develop into a full outburst. By May 17.25, the object was back to quiescence. The peak amplitude of the outburst was 0.34 ± 0.02 mag brighter than the quiescent mean in R and occurred on May 12.37. Two observations with the PCA instrument on RXTE lasting 1500 and 2500 s were taken on May 14.38 and 16.91, respectively. There is an indication that the source was detected with the PCA, but the count rates were heavily background-dominated and low enough that preliminary analysis does not allow a definitive determination of the flux. We therefore quote an upper limit of 4 × 10^{−11} erg cm^{−2} s^{−1} (2.5–20 keV). A similar truncated ‘minioutburst’ also occurred in Sept. 1998 (IAUC 7017, 7025).”

SUPERNOVA 2000ce IN UGC 4195 K. Krisciunas and A. Rest, University of Washington, report prelimi-

nary photometry obtained by Rest, R. McMillan, and N. C. Hastings using the Apache Point Observatory 3.5-m telescope (and software aperture of diameter 4′′.5): May 11.12 UT, V = 16.92, B − V = +0.74, V − R = +0.35, V −I = +0.33 (uncertainties ± 0.03 mag). On May 20.12, V = 17.45 ± 0.03, B−V = +1.15 ± 0.07, V −R = +0.39 ± 0.04, V −I = +0.63 ± 0.04 (software aperture of diameter 5′′.6). On May 11.25, they found infrared magnitudes J = 16.65 ± 0.07, H = 15.85 ± 0.07, K = 15.99 ± 0.10 (software aperture of diameter 5′′.7; sky annulus of radius 5′′.7–9′′.5). They add: “If the estimated time of maximum light (IAUC 7422) of May 4 is correct and this object’s V - minus-near-infrared color evolution is similar to the supernovae studied by

Krisciunas et al. (1999, http://xxx.lanl.gov/abs/astro-ph/?9912219),

then the extinction along the line-of-sight to this object is A_{V} ∼ 2.3 mag, which would explain the red B − V color so soon after maximum.”

Additional CCD R magnitude estimates by K. Hornoch, Lelekovice, Czech Republic (cf. IAUC 7424): May 12.837 UT, 16.2; 13.845, 16.1; 14.846, 16.2; 15.848, 16.3; 23.857, 16.2.

2000 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green